### v1

@@ 4.6. Threats to Validity
Quantum data encoding is a major threat to validity. Current classical computers can only simulate a limited number of qubits due to the enormous computational cost incurred by numerical simulations of high-dimensional quantum systems. Classical data is typically downsampled and then encoded into quantum states for training QNNs. We alleviate this threat by employing the Fashion and MNIST datasets, characterized by lower data dimensions.
Another threat could be the selection of QNN models. The construction of QNNs is still under exploration, and there exist various QNN variants based on different underlying architectures. To ensure the generalizability of QuanTest, we employ three different architectures of QNN models to evaluate QuanTest’s performance. We believe that QuanTest can be applied to the majority of PQC-based QNNs.
@@ 5.2. Adversarial Machine Learning
In recent years, vulnerabilities to adversarial examples have also been found in quantum machine learning models (Lu et al., 2020; Gong and Deng, 2022). Quantum adversarial learning has been experimentally demonstrated on programmable superconducting qubits (Ren et al., 2022). In contrast to these works, QuanTest systematically tests QNN systems by entanglement guidance. Moreover, QuanTest is exceptionally scalable and applicable to various PQC-based QNN systems and quantum state encoding strategies. Modularity allows QuanTest to directly achieve most quantum-adapted adversarial attack methods without the need for additional adjustments and adaptations, thus generating quantum adversarial examples with high efficiency.

### v2

@@ 5.5. Threats to Validity
Quantum data encoding is a major threat to validity. Current classical computers can only simulate a limited number of qubits due to the enormous computational cost incurred by numerical simulations of high-dimensional quantum systems (Biamonte et al., 2017). Classical data is typically downsampled and then encoded into quantum states for training QNNs (Hur et al., 2022). We alleviate this threat by using two popular datasets with lower data dimensions Fashion and MNIST, for evaluation.
Another threat could be the selection of QNN models. The construction of QNNs is still under exploration, and there exist various QNN variants based on different underlying architectures (Steinbrecher et al., 2019; Killoran et al., 2019). To ensure the generalizability of QuanTest, we employ three different architectures of QNN models to evaluate QuanTest’s performance. We believe that QuanTest can be applied to the majority of PQC-based QNNs, and the testing of other QNN systems with different architectures is worth exploring in future work.
Finally, variations in quantum computer implementation technologies also pose a potential threat. To guarantee the applicability of QuanTest in the current NISQ era, all methods we use are based on demonstrations on real devices. Specifically, fidelity, as a similarity evaluation metric, can be obtained by performing the SWAP test. For the core part of QuanTest, Ren et al. (Ren et al., 2022) demonstrated the experimental generation of quantum adversarial examples using an array of ten programmable superconducting transmon qubits. Similarly, Haug et al. (Haug and Kim, 2023) demonstrated the Bell measurement protocol for MW entanglement measurement on the IonQ quantum computer.
@@ 6.2. Adversarial Machine Learning
In recent years, vulnerabilities to adversarial examples have also been found in quantum machine learning models (Lu et al., 2020; Gong and Deng, 2022). Quantum adversarial learning has been experimentally demonstrated on programmable superconducting qubits (Ren et al., 2022). In contrast to these works, QuanTest systematically tests QNN systems by entanglement guidance. Moreover, QuanTest is exceptionally scalable and applicable to various PQC-based QNN systems and quantum state encoding strategies. Modularity allows QuanTest to directly achieve most quantum-adapted adversarial attack methods without the need for additional adjustments and adaptations, thus generating various quantum adversarial examples with high efficiency.
@@ 7. Conclusion
Acknowledgements.
We thank Jia Liao and Yuhu Lu for assistance with the writing and revision process.